# Deterministic Dynamic Edge-Colouring

### v1

@@ 3 Preliminaries & Notation
The first parts of this section is very similar to Section 2 of the full-version of [Chr23], as we need similar terminology. For a positive integer $t$, we let $[t]=\{1,2,\dots,t\}$ be the set of numbers in $\mathbb{N}$ that are greater than 00, but at most $t$. We let $G=(V,E)$ be a graph on $n$ vertices and $m$ edges. A subgraph $H\subset G$ of $G$ is a graph such that $V(H)\subset V(G)$ and $E(H)\subset E(G)$. The $t$-hop neighbourhood of $H$ in $G$, $N^{t}(H)$ consists of all of the vertices in $G$ of distance at most $t$ (in $G$) to a vertex in $H$. In particular, $H\subset N(H)$. A proper (partial) $k$-edge-colouring $c$ of $G$ is a function $c:E(G)\mapsto[k]\cup\{\neg\}$ such that if $c(e)\neq\neg$ and $c(e^{\prime})\neq\neg$, then $e\cap e^{\prime}$ implies that $c(e)\neq c(e^{\prime})$. An edge $e$ such that $c(e)=\neg$ is said to be uncoloured. Two uncoloured edges may share an endpoint in a proper partial colouring. If no edges are uncoloured, we say that the colouring is a proper $k$-edge-colouring. For a vertex set $U\subset V(G)$ (resp. edge set $U^{\prime}\subset E(G)$), we let $G[U]$ (resp. $G[U^{\prime}]$) be the induced graph of $U$ in $G$ i.e. the subgraph of $G$ on vertex set $V(G[U])=U$ containing the maximum number of edges (resp. the induced graph of $U^{\prime}$ in $G$ i.e. the subgraph with edge set $E(G[U^{\prime}])=U^{\prime}$ and vertex set $V(G[U^{\prime}])=V(G)$). Given a proper (partial) $k$-edge-colouring $c$ and a set of colours $S\subset[k]$, we let $G[S]$ be the graph containing all edges coloured with a colour from $S$, and all of the vertices incident to an edge coloured with a colour from $S$. Given a proper (partial) $t$-edge-colouring of a graph $G$, we define the set $A(v)$ of available colours at a vertex $v$ to consist of precisely the colours of $[t]$ that no edge incident to $v$ has received.
For a directed graph $G$, any edge denoted by $(uv)\in E(G)$ is oriented from $u$ towards $v$. We denote by $d_{\text{in}}(v)$ (resp. $d_{\text{out}}(v)$) the in-degree of $v$ (resp. out-degree of $v$). Similarly, if $E^{\prime}\subset E(G)$ is a subset of edges and any subgraph $H$, we let $N^{j}_{E^{\prime},\text{in}}(H)$ denote the $j$-hop in-neighbourhood of $H$ in the graph induced by $E^{\prime}$. Similarly, we let $E_{E^{\prime},\text{in}}(H)$ be the edge set containing all in-edges of $H$ also belonging to $E^{\prime}$.
@@ 3.1 Chains and shifts
In order to extend a partial edge-colouring, we will have to assign some coloured edge new colours. In order to describe such a process formally, we briefly recall the chain and shift terminology as it was used by for example Bernshteyn [Ber22] (see [Ber22] for a more in-depth treatment).
@@
$\operatorname{Shift}(c,e_{1},e_{2})(e)=\begin{cases}c(e_{2})&\quad\text{if }e=e_{1}\\ \neg&\quad\text{if }e=e_{2}\\ c(e)&\quad\text{if }e\notin\{e_{1},e_{2}\}\\ \end{cases}$
@@
A chain $C$ of size $k$ is then a set of edges $C=(e_{1},\dots,e_{k})$ such that $e_{i}$ and $e_{i+1}$ are adjacent for all $i$ and $c(e_{i})=\neg$ if and only if $i=1$. For $10\leq j\leq k-$ 1, we can $j$-shift such a chain by performing $\operatorname{Shift}_{j}(c,C)$ defined as:
@@
The initial segment of length $s$ of a chain $C$ is then the chain $C|s=(e_{1},\dots,e_{s})$. One can concatenate two chains $C_{1}=(e_{1},\dots,e_{s_{1}})$ and $C_{2}=(e_{s_{1}},f_{2},\dots,f_{s_{2}})$ to get the chain $C_{1}+C_{2}=(e_{1},\dots,e_{s_{1}},f_{2},\dots,f_{s_{2}})$. Then in order to shift such a chain, we use $\operatorname{Shift}(c,C_{1}+C_{2})=\operatorname{Shift}(\operatorname{Shift}(c,C_{1}),C_{2})$.
A fan chain with center $u$ is a chain $F$ of the form $F=(uw_{1},\dots,uw_{k})$. Here $u$ is the center of the fan, denoted by $\text{center}(F)$. Bernshteyn [Ber22] showed that for such fans to be shiftable chains, we require $c(uw_{i+1})\in A(w_{i})$. We will let the colour of $c(uw_{i+1})$ be the representative available colour at $w_{i}$. We will also pick a representative available colour at $w_{k}$ in $A(w_{k})$. If this representative available colour is also either available at $u$ or if it is the representative available colour for some $w_{j}$ with $j<k$, then we say that the fan is maximal.
@@
Vizing originally showed how to choose both a fan chain $F$ and a bichromatic path chain $P$ that together form a bigger chain $F+P$ such that $\operatorname{Shift}(c,F+P)$ leaves some edge $uv$ with an available colour. This type of chain has been referred to as a Vizing chain in the literature [Ber22, BD23, Chr23, GP20].
Definition 9.
[…]
@@
A (multi-step) Vizing chain is augmenting if shifting the chain leaves the final edge of the chain with a colour that is available at both endpoints of the edge. If the chain is not augmenting, it is truncated. The length of a multi-step Vizing chain is the sum of the lengths of the chains $F_{j}+P_{j}$. We will only be considering Vizing chains that are non-overlapping:
Definition 11.
An $i$-step Vizing chain is non-overlapping if every pair of Vizing chains $F_{j}+P_{j}$ on edge $e_{j}$ and $F_{k}+P_{k}$ on edge $e_{k}$ share an edge exactly when $j=k-1$ and the shared edge is $e_{k}$, and if, furthermore, for all $k\leq i$ no edge in the chain $F_{k}+P_{k}=(e_{1},\dots,e_{s})$ is repeated i.e. $e_{j}=e_{h}$ iff $j=h$.
@@ 3.2 Algorithms for constructing and extending Vizing chains
In this subsection, we briefly recall some well-known algorithms for constructing Vizing chains. In some cases, one is allowed to make arbitrary choices in the constructions. In order to perform the constructions in a consistent manner, we will order the colours $[s]$ according to the strict ordering $<$ on $\mathbb{N}$. Now, whenever an arbitrary choice between colours (possibly satisfying some conditions) can be made, we will always choose the smallest colour according to $<$. We say that a fan is consistent if the representative available colour $\alpha_{i}$ of each edge $uw_{i}$ is the smallest colour in $A(w_{i})$ subject to the requirements that 1)1)1 ) if $A(u)\cap A(w_{i})\neq\emptyset$, then $\alpha_{i}\in A(u)\cap A(w_{i})$ and 2)2)2 ) else if $j$ is the smallest integer $j<i$ with $\alpha_{j}\in A(w_{i})$ then $\alpha_{i}=\alpha_{j}$. Given a consistent fan-chain $F=(uw_{1},\dots,uw_{k})$, a $(\kappa_{1},\kappa_{2})$-bichromatic path chain $P$ is consistent with $F$ if: $\kappa_{1}\in A(u)$, $\kappa_{2}\in A(w_{k})$ are also chosen as small as possible according to $<$.
@@
Lemma 12 ([Ber22, BD23, Viz64]).
@@
    else if there exists a $h<i$ such that $\alpha_{h}\in A(w_{j})$, we let $h^{\prime}$ be the smallest such $h$ and set $\alpha_{j}=\alpha_{h}$. In this case, the construction terminates, and we conclude that $F$ is unique by the induction hypothesis.
@@
The above construction must terminate. Indeed, when $j=d(u)-1$ the third case cannot occur, as $u$ has at most $d(u)-1$ coloured edges incident to it.
As for the furthermore part, we can check 1)1)1 ) in $\tilde{O}(\Delta_{\max}^{2})$ by just checking for each colour in $A(w_{j})$ if it belongs to $A(u)$, we can check 2)2)2 ) in $O(\Delta_{\max}^{2})$ time, since we can check if $\alpha_{k}\in A(w_{j})$ in $O(\Delta_{\max})$ time. Finally, we can do 3)3)3 ) in $\tilde{O}(\Delta_{\max})$ time by simply computing $A(w_{j})$ and extracting the minimum. ∎
Given a consistent maximal fan chain $F=(uw_{1},\dots,uw_{k})$ and a parameter $t\in\mathbb{Z}_{\geq 0}$, we can extend $F$ to a Vizing chain $(F^{\prime}+P)_{t}$, or simply $F^{\prime}+P$ if $t$ is clear from the context, in a consistent manner (note $F^{\prime}$ might not be equal to $F$), as follows: if $\alpha_{k}\in A(u)$, we set $P=(uw_{k})$ and $F^{\prime}=F$. Otherwise, we let $\kappa_{1}$ be the smallest colour in $A(u)$ and $\kappa_{2}=\alpha_{k}$. Let $P^{\prime}=p_{1},p_{2},\dots,p_{t^{\prime}}$ be the unique maximal $(\kappa_{1},\kappa_{2})$-bichromatic path with $p_{1}=w_{k}$. If $t<t^{\prime}-1$, we let $F^{\prime}=F$, and we let $P=(uw_{k},w_{k}p_{2},\dots p_{t}p_{t+1})$. Otherwise, if $t\geq t^{\prime}-1$, there are $3$ cases:
@@
if $p_{t^{\prime}-1}=w_{i+1}$ and $p_{t^{\prime}}=u$ with $i+1<k$, we let $F^{\prime}=(uw_{1},\dots,uw_{i+1})$ and we let $P=(uw_{i+1},w_{i+1}p_{t^{\prime}-2},p_{t^{\prime}-2}p_{t^{\prime}-3},\dots,p_{2}p_{1})$. Note that in this case $F^{\prime}$ is not a maximal fan.
@@
It is straight-forward to check that we can construct $F^{\prime}+P$ in $O(\Delta_{\max}^{3}+\Delta_{\max}\cdot{}t)$ time, and that the construction is unique. If $e=uv$ is any uncoloured edge and $F$ is the unique maximal fan chain guaranteed by Lemma 12 centered on $u$, then for any given $t$, we say that $(F^{\prime}+P)_{t}$ is a consistent Vizing chain with center $u$. Note that $(F^{\prime}+P)_{t}$ is in fact unique.
Next we will see a different algorithm needed to extend any $i$-step Vizing chain to an $(i+1)$-step Vizing chain in a consistent way. Again we will first construct a fan chain, and then explain how to compute the accompanying bichromatic path chain. We will use the so-called ’second fan lemma’ [Ber22, GP20].
Lemma 13 ([Ber22, GP20]).
Let $G$ be a graph with maximum degree $\Delta$ satisfying $\Delta\leq\Delta_{\max}$. Let $c$ be a proper partial $(\Delta_{\max}+1)$-edge-colouring, and let $e=uv\in E(G)$ be an uncoloured edge. Let $\kappa_{1}\in[\Delta+1]$ be an available colour at $u$ and let $\kappa_{2}\in[\Delta+1]$ be an available colour at $v$. Then there exists a fan chain $uw_{1},\dots,uw_{k}$ with $w_{1}=v$ such that either
@@
If the representative available colours are chosen as small as possible subject to first 1)1)1 ), then 2)2)2 ), and finally 3)3)3 ) above, then $F$ is unique.
Furthermore, we may compute $F$ in $\tilde{O}(\Delta_{\max}^{3})$ time.
Proof.
We construct a maximal fan almost exactly as in the proof of Lemma 12. Again we begin by setting $w_{1}=v$. We pick an available representative colour $\alpha_{1}$ in $A(v)$ for $v$ as follows: if $A(v)\cap A(u)\neq\emptyset$, we let $\alpha_{1}$ be the smallest element of $A(v)\cap A(u)$. In this case, the construction terminates, and we conclude that $F$ is unique. Otherwise, we let $\alpha_{1}$ be the smallest element of $A(v)\setminus\{\kappa_{2}\}$. Note that since $uv$ is uncoloured, $|A(v)|\geq 2$. We let $w_{2}$ be the unique neighbour of $u$ with $c(uw_{2})=\alpha_{1}$. In this case, we note that any consistent maximal fan chain must have $(uw_{1},uw_{2})$ as an initial segment.
@@
    else if there exists a $h<i$ such that $\alpha_{h}\in A(w_{j})$, we let $h^{\prime}$ be the smallest such $h$ and set $\alpha_{j}=\alpha_{h}$. In this case, the construction terminates, and we conclude that $F$ is unique by the induction hypothesis.
@@
The above construction must terminate. Indeed, when $j=d(u)-1$ the third case cannot occur, as $u$ has at most $d(u)-1$ coloured edges incident to it.
Similarly to Lemma 12, we can construct the fan in $\tilde{O}(\Delta_{\max}^{3})$ time. ∎
@@
Now given some $i$-step Vizing chain $F_{1}+P_{1}+F_{2}+P_{2}+\dots F_{i}+P_{i}$ and some parameter $t\in\mathbb{Z}_{\geq 0}$, we can extend it in a consistent manner by applying Lemma 13 to the colouring $\operatorname{Shift}(c,F_{1}+P_{1}+F_{2}+P_{2}+\dots F_{i}+P_{i})$ and the final edge $p_{t_{i}-1}p_{t_{i}}=u^{\prime}v^{\prime}$ to get a unique fan $F=(u^{\prime}w_{1}^{\prime},u^{\prime}w_{2}^{\prime},\dots,u^{\prime}w_{k}^{\prime})$. Let $P_{i}$ be $(\kappa_{1},\kappa_{2})$-bichromatic with $c(u^{\prime}v^{\prime})=\kappa_{2}$, then we apply Lemma 13 with $\kappa_{1}$ and $\kappa_{2}$. If we are in Case $1$, we set $F_{i+1}=F$ and $P_{i+1}=u^{\prime}w_{k}^{\prime}$. If we are in Case $2$, we set $F_{i+1}=F$ and let $\tau_{1}\in A(u^{\prime})\setminus\{\kappa_{1},\kappa_{2}\}$ and $\tau_{2}=\alpha_{k}$ be the smallest available colours. Note that $|A(u^{\prime})|\geq 2$ in the shifted colouring, and that $\kappa_{2}\notin A(u)$ by assumption. Then we let $P^{\prime}=p^{\prime}_{1}p^{\prime}_{2}\dots p^{\prime}_{t^{\prime}}$ be the maximal $(\tau_{1},\tau_{2})$-bichromatic path beginning at $p^{\prime}_{1}=w_{k}^{\prime}$. If $t<t^{\prime}-1$, we set $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t}p_{t+1})$, otherwise we set $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t^{\prime}-1}p_{t^{\prime}})$. Otherwise, we are in Case 3)3)3 ). Again we set $F_{i+1}=F$, but this time we let $P^{\prime}=p^{\prime}_{1}p^{\prime}_{2}\dots p^{\prime}_{t^{\prime}}$ be the maximal $(\kappa_{1},\kappa_{2})$-bichromatic path beginning at $p^{\prime}_{1}=w_{k}^{\prime}$. If $t<t^{\prime}-1$, we set $P_{i+1}=(u^{\prime}w_{k}^{\prime},w^{\prime}_{k}p_{2},\dots p_{t}p_{t+1})$, otherwise we set $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t^{\prime}-1}p_{t^{\prime}})$.

### v2

@@ 3 Preliminaries & Notation
The first parts of this section is very similar to Section 2 of the full-version of [Christiansen], as we need similar terminology. For a positive integer $t$, we let $t]=\{1,2,\dots,t\}$ be the set of numbers in $\mathbb{N}$ that are greater than 0, but at most $t$. We let $G=(V,E)$ be a graph on $n$ vertices and $m$ edges. A subgraph $H\subset G$ of $G$ is a graph such that $V(H)\subset V(G)$ and $E(H)\subset E(G)$. The $t$-hop neighbourhood of $H$ in $G$, $N^{t}(H)$ consists of all of the vertices in $G$ of distance at most $t$ (in $G$) to a vertex in $H$. In particular, $H\subset N(H)$. A proper (partial) $k$-edge-colouring $c$ of $G$ is a function $c:E(G)\mapsto[k]\cup\{\neg\}$ such that if $c(e)\neq\neg$ and $c(e^{\prime})\neq\neg$, then $e\cap e^{\prime}$ implies that $c(e)\neq c(e^{\prime})$. An edge $e$ such that $c(e)=\neg$ is said to be uncoloured. Two uncoloured edges may share an endpoint in a proper partial colouring. If no edges are uncoloured, we say that the colouring is a proper $k$-edge-colouring. For a vertex set $U\subset V(G)$ (resp. edge set $U^{\prime}\subset E(G)$), we let $G[U]$ (resp. $G[U^{\prime}]$) be the induced graph of $U$ in $G$ i.e. the subgraph of $G$ on vertex set $V(G[U])=U$ containing the maximum number of edges (resp. the induced graph of $U^{\prime}$ in $G$ i.e. the subgraph with edge set $E(G[U^{\prime}])=U^{\prime}$ and vertex set $V(G[U^{\prime}])=V(G)$). Given a proper (partial) $k$-edge-colouring $c$ and a set of colours $S\subset[k]$, we let $G[S]$ be the graph containing all edges coloured with a colour from $S$, and all of the vertices incident to an edge coloured with a colour from $S$. Given a proper (partial) $t$-edge-colouring of a graph $G$, we define the set $A(v)$ of available colours at a vertex $v$ to consist of precisely the colours of $[t]$ that no edge incident to $v$ has received.
For a directed graph $G$, any edge denoted by $(uv)\in E(G)$ is oriented from $u$ towards $v$. We denote by $d_{\text{in}}(v)$ (resp. $d_{\text{out}}(v)$) the in-degree of $v$ (resp. out-degree of $v$). Similarly, if $E^{\prime}\subset E(G)$ is a subset of edges and $H$ any subgraph, we let $N^{j}_{E^{\prime},\text{in}}(H)$ denote the $j$-hop in-neighbourhood of $H$ in the graph induced by $E^{\prime}$. Similarly, we let $E_{E^{\prime},\text{in}}(H)$ be the edge set containing all in-edges of $H$ also belonging to $E^{\prime}$.
@@ 3.1 Chains and shifts
In order to extend a partial edge-colouring, we will have to assign some coloured edges new colours. In order to describe such a process formally, we briefly recall the chain and shift terminology as it was used by for example Bernshteyn [2BERNSHTEYN] (see [2BERNSHTEYN] for a more in-depth treatment).
@@
[…]
@@
A chain $C$ of size (or length) $k$ is then a set of edges $C=(e_{1},\dots,e_{k})$ such that $e_{i}$ and $e_{i+1}$ are adjacent for all $i$ and $c(e_{i})=\neg$ if and only if $i=1$. For $0\leq j\leq k-1$, we can $j$-shift such a chain by performing $\operatorname{Shift}_{j}(c,C)$ defined as:
@@
The initial segment of length $s$ of a chain $C$ is then the chain $C|s=(e_{1},\dots,e_{s})$. One can concatenate two chains $C_{1}=(e_{1},\dots,e_{s_{1}})$ and $C_{2}=(e_{s_{1}},f_{2},\dots,f_{s_{2}})$ to get the chain $C_{1}+C_{2}=(e_{1},\dots,e_{s_{1}},f_{2},\dots,f_{s_{2}})$. The chain $C_{1}+C_{2}$ is $c$-shiftable if $C_{1}$ is $c$-shiftable and $C_{2}$ is $\operatorname{Shift}(c,C_{1})$-shiftable. In order to shift such a chain, we use $\operatorname{Shift}(c,C_{1}+C_{2})=\operatorname{Shift}(\operatorname{Shift}(c,C_{1}),C_{2})$.
A fan chain with center $u$ is a chain $F$ of the form $F=(uw_{1},\dots,uw_{k})$. Here $u$ is the center of the fan, denoted by $\text{center}(F)$. Bernshteyn [2BERNSHTEYN] showed that for such fans to be shiftable chains, we require $c(uw_{i+1})\in A(w_{i})$. We will let the colour of $c(uw_{i+1})$ be the representative available colour at $w_{i}$. We will also pick a representative available colour at $w_{k}$ in $A(w_{k})$. If this representative available colour is also either available at $u$ or if it is the representative available colour for some $w_{j}$ with $j<k$, then we say that the fan is maximal.
@@
Vizing originally showed how to choose both a fan chain $F$ and a bichromatic path chain $P$ that together form a bigger chain $F+P$ such that $\operatorname{Shift}(c,F+P)$ leaves some edge $uv$ with an available colour. This type of chain has been referred to as a Vizing chain in the literature [2BERNSHTEYN, BernDhawan, Christiansen, grebik2020measurable].
Definition 3.1.
Given a proper partial colouring $c$ of $G$ and an edge $e=uv\in E(G)$ with $c(e)=\neg$, a Vizing chain on $e$ with center $u$ is a chain $F+P$, where $F=(uw_{1},\dots,uw_{k})$ is a fan chain with $\text{center}(F)=u$ and $P=(p_{0}p_{1},\dots,p_{t-1}p_{t})$ is a $(\kappa_{1},\kappa_{2})$-bichromatic path chain with $p_{0}=u$, $p_{1}=w_{k}$, $\kappa_{1}\in A(u)$, and $\kappa_{2}\in A(w_{k})$.
@@
The length of a multi-step Vizing chain is the sum of the lengths of the chains $F_{j}+P_{j}$. We will only be considering Vizing chains that are non-overlapping:
Definition 3.3.
An $i$-step Vizing chain is non-overlapping if every pair of Vizing chains $F_{j}+P_{j}$ on edge $e_{j}$ and $F_{k}+P_{k}$ on edge $e_{k}$ share an edge exactly when j=k1j=k-1 and the shared edge is $e_{k}$, and if, furthermore, for all $k\leq i$ no edge in the chain $F_{k}+P_{k}=(e_{1},\dots,e_{s})$ is repeated i.e. $e_{j}=e_{h}$ iff $j=h$.
@@ 3.2 Algorithms for constructing and extending Vizing chains
In this subsection, we briefly recall some well-known algorithms for constructing Vizing chains. In some cases, one is allowed to make arbitrary choices in the constructions. In order to perform the constructions in a consistent manner, we will order the colours $[s]$ according to the strict ordering << on $\mathbb{N}$. Now, whenever an arbitrary choice between colours (possibly satisfying some conditions) can be made, we will always choose the smallest colour according to <<. We say that a fan is consistent if the representative available colour $\alpha_{i}$ of each vertex $w_{i}$ is the smallest colour in $A(w_{i})$ subject to the requirements that $1)$ if $A(u)\cap A(w_{i})\neq\emptyset$, then $\alpha_{i}\in A(u)\cap A(w_{i})$ and $2)$ else if $j$ is the smallest integer $j<i$ with $\alpha_{j}\in A(w_{i})$ then $\alpha_{i}=\alpha_{j}$. Given a consistent fan-chain $F=(uw_{1},\dots,uw_{k})$, a $(\kappa_{1},\kappa_{2})$-bichromatic path chain $P$ is consistent with $F$ if: $\kappa_{1}\in A(u)$ is chosen as small as possible according to << and $\kappa_{2}=\alpha_{k}$.
@@
Lemma 3.4 ([2BERNSHTEYN, BernDhawan, vizing1964estimate]).
@@
    else if there exists a $h<i$ such that $\alpha_{h}\in A(w_{j})$, we let $h^{\prime}$ be the smallest such $h$ and set $\alpha_{j}=\alpha_{h^{\prime}}$. In this case, the construction terminates, and we conclude that $F$ is unique by the induction hypothesis.
@@
The above construction must terminate. Indeed, when j=d(u)1j=d(u)-1 the third case cannot occur, as $u$ has at most d(u)1d(u)-1 coloured edges incident to it.
As for the furthermore part, we can check $1)$ in $\tilde{O}(\Delta_{\max})$ by just checking for each colour in $A(w_{j})$ if it belongs to $A(u)$, we can check $2)$ in $O(\Delta_{\max}^{2})$ time, since we can check if $\alpha_{k}\in A(w_{j})$ in $O(\Delta_{\max})$ time. Finally, we can do $3)$ in $\tilde{O}(\Delta_{\max})$ time by simply computing $A(w_{j})$ and extracting the minimum.
Given a consistent maximal fan chain $F=(uw_{1},\dots,uw_{k})$ and a parameter $t\in\mathbb{Z}_{\geq 0}$, we can extend $F$ to a Vizing chain $(F^{\prime}+P)_{t}$, or simply $F^{\prime}+P$ if $t$ is clear from the context, in a consistent manner (note $F^{\prime}$ might not be equal to $F$), as follows: if $\alpha_{k}\in A(u)$, we set $P=(uw_{k})$ and $F^{\prime}=F$. Otherwise, we let $\kappa_{1}$ be the smallest colour in $A(u)$ and $\kappa_{2}=\alpha_{k}$. Let $P^{\prime}=p_{1},p_{2},\dots,p_{t^{\prime}}$ be the unique maximal $(\kappa_{1},\kappa_{2})$-bichromatic path with $p_{1}=w_{k}$. If $t<t^{\prime}-1$, we let $F^{\prime}=F$, and we let $P=(uw_{k},w_{k}p_{2},\dots p_{t}p_{t+1})$ (except for the case $t=t^{\prime}-2$ in Case 2 below, where we do something else, as explained later). Otherwise, if $t\geq t^{\prime}-1$, there are $3$ cases:
@@
if $p_{t^{\prime}-1}=w_{i+1}$ and $p_{t^{\prime}}=u$ with $i+1<k$, we let $F^{\prime}=(uw_{1},\dots,uw_{i+1})$ and we let $P=(uw_{i+1},w_{i+1}p_{t^{\prime}-2},p_{t^{\prime}-2}p_{t^{\prime}-3},\dots,p_{2}p_{1})$. Note that in this case $F^{\prime}$ is not a maximal fan. In the case where $t=t^{\prime}-2$, we pick $F^{\prime}=(uw_{1},\dots,uw_{i+1})$ and we let $P=(uw_{i+1},w_{i+1}p_{t^{\prime}-2},p_{t^{\prime}-2}p_{t^{\prime}-3},\dots,p_{3}p_{2})$
@@
It is straight-forward to check that we can construct $F^{\prime}+P$ in $\tilde{O}(\Delta_{\max}^{3}+\Delta_{\max}\cdot{}t)$ time, and that the construction is unique. If $e=uv$ is any uncoloured edge and $F$ is the unique maximal fan chain guaranteed by Lemma 3.4 centered on $u$, then for any given $t$, we say that $(F^{\prime}+P)_{t}$ is a consistent Vizing chain with center $u$. Note that $(F^{\prime}+P)_{t}$ is unique.
Next we will see a different algorithm needed to extend any $i$-step Vizing chain to an $(i+1)$-step Vizing chain in a consistent way. Again we will first construct a fan chain, and then explain how to compute the accompanying bichromatic path chain. We will use the so-called ’second fan lemma’ [2BERNSHTEYN, grebik2020measurable].
Lemma 3.5 ([2BERNSHTEYN, grebik2020measurable]).
Let $G$ be a graph with maximum degree $\Delta$ satisfying $\Delta\leq\Delta_{\max}$. Let $c$ be a proper partial $(\Delta_{\max}+1)$-edge-colouring, and let $e=uv\in E(G)$ be an uncoloured edge. Let $\kappa_{1}\in[\Delta_{\max}+1]$ be an available colour at $u$ and let $\kappa_{2}\in[\Delta_{\max}+1]$ be an available colour at $v$. Then there exists a fan chain $uw_{1},\dots,uw_{k}$ with $w_{1}=v$ such that either
@@
If the representative available colours are chosen as small as possible subject to first $1)$, then $2)$, and finally $3)$ above, then $F$ is unique.
Furthermore, we can compute $F$ in $\tilde{O}(\Delta_{\max}^{3})$ time.
Proof 2.
We construct a maximal fan almost exactly as in the proof of Lemma 3.4. Again we begin by setting $w_{1}=v$. We pick an available representative colour $\alpha_{1}$ in $A(v)$ for $v$ as follows: if $A(v)\cap A(u)\neq\emptyset$, we let $\alpha_{1}$ be the smallest element of $A(v)\cap A(u)$. In this case, the construction terminates, and we conclude that $F$ is unique. Otherwise, we let $\alpha_{1}$ be the smallest element of $A(v)\setminus\{\kappa_{2}\}$. Note that since $uv$ is uncoloured, $|A(v)|\geq 2$. We let $w_{2}$ be the unique neighbour of $u$ with $c(uw_{2})=\alpha_{1}$. In this case, we note that any consistent maximal fan chain must have $(uw_{1},uw_{2})$ as an initial segment.
@@
    else if there exists a $h<i$ such that $\alpha_{h}\in A(w_{j})$, we let $h^{\prime}$ be the smallest such $h$ and set $\alpha_{j}=\alpha_{h^{\prime}}$. In this case, the construction terminates, and we conclude that $F$ is unique by the induction hypothesis.
@@
The above construction must terminate. Indeed, when j=d(u)1j=d(u)-1 the third case cannot occur, as $u$ has at most d(u)1d(u)-1 coloured edges incident to it.
Similarly to Lemma 3.4, we can construct the fan in $\tilde{O}(\Delta_{\max}^{3})$ time.
@@
Now given some $i$-step Vizing chain $F_{1}+P_{1}+F_{2}+P_{2}+\dots F_{i}+P_{i}$ and some parameter $t\in\mathbb{Z}_{\geq 0}$, we can extend it in a consistent manner by applying Lemma 3.5 to the colouring $\operatorname{Shift}(c,F_{1}+P_{1}+F_{2}+P_{2}+\dots F_{i}+P_{i})$ and the final edge $p_{t_{i}-1}p_{t_{i}}=u^{\prime}v^{\prime}$ to get a unique fan $F=(u^{\prime}w_{1}^{\prime},u^{\prime}w_{2}^{\prime},\dots,u^{\prime}w_{k}^{\prime})$. Let $P_{i}$ be $(\kappa_{1},\kappa_{2})$-bichromatic with $c(u^{\prime}v^{\prime})=\kappa_{2}$, then we apply Lemma 3.5 with $\kappa_{1}$ and $\kappa_{2}$. If we are in Case $1.$, we set $F_{i+1}=F$ and $P_{i+1}=u^{\prime}w_{k}^{\prime}$. If we are in Case $2.$, we let $\tau_{1}\in A(u^{\prime})\setminus\{\kappa_{1},\kappa_{2}\}$ be the smallest available colour and $\tau_{2}=\alpha_{k}$. Note that $|A(u^{\prime})|\geq 2$ in the shifted colouring, and that $\kappa_{2}\notin A(u)$ by assumption. Then we let $P^{\prime}=p^{\prime}_{1}p^{\prime}_{2}\dots p^{\prime}_{t^{\prime}}$ be the maximal $(\tau_{1},\tau_{2})$-bichromatic path beginning at $p^{\prime}_{1}=w_{k}^{\prime}$.
If $p^{\prime}_{t}=u$ and $p^{\prime}_{t-1}=w_{s}^{\prime}$ for some $s$, there are three sub-cases. If $t<t^{\prime}-2$, we set $F_{i+1}=F$ and $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t-1}p_{t})$. Else if $t=t^{\prime}-1$, we set $F=(uw_{1}^{\prime},\dots,uw_{s}^{\prime})$ and $P_{i+1}=(u^{\prime}w_{s}^{\prime},w_{s}^{\prime}p^{\prime}_{t^{\prime}-2},\dots p_{3}p_{2})$. If none of the above hold, it must be the case that $t>t^{\prime}-1$, and then we set $F=(uw_{1}^{\prime},\dots,uw_{s}^{\prime})$ and $P_{i+1}=(u^{\prime}w_{s}^{\prime},w_{s}^{\prime}p^{\prime}_{t^{\prime}-2},\dots p_{2}p_{1})$.
If $p^{\prime}_{t}=w_{s}^{\prime}$ for some $s$, there are two sub-cases. If $t<t^{\prime}-1$, we set $F_{i+1}=F$ and $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t-1}p_{t})$. Else, we set $F=(uw_{1}^{\prime},\dots,uw_{s}^{\prime})$ and $P_{i+1}=(u^{\prime}w_{s}^{\prime},w_{s}^{\prime}p^{\prime}_{t^{\prime}-1},\dots p_{2}p_{1})$.
Else $p^{\prime}_{t}$ is disjoint from $F-\{w^{\prime}_{k}\}$, in which case there are $2$ sub-cases: If $t<t^{\prime}-1$, we set $F_{i+1}=F$ and $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t-1}p_{t})$. Else, we set $F_{i+1}=F$ and $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t^{\prime}-1}p_{t^{\prime}})$.
Otherwise, we are in Case $3)$. Again we set $F_{i+1}=F$, but this time we let $P^{\prime}=p^{\prime}_{1}p^{\prime}_{2}\dots p^{\prime}_{t^{\prime}}$ be the maximal $(\kappa_{1},\kappa_{2})$-bichromatic path beginning at $p^{\prime}_{1}=w_{k}^{\prime}$. If $t<t^{\prime}-1$, we set $P_{i+1}=(u^{\prime}w_{k}^{\prime},w^{\prime}_{k}p_{2},\dots p_{t}p_{t+1})$, otherwise we set $P_{i+1}=(u^{\prime}w_{k}^{\prime},w_{k}p^{\prime}_{2},\dots p_{t^{\prime}-1}p_{t^{\prime}})$.